I need your help. I am not the greatest when it comes to solving functions.
x^2=1+y^2....is this a function or not
is this correct?
y=square root x-1
x-1>0 or x>1
domain [1,infin)
range [0,infin)
no function
Is this correct?
y=2x-3
2(1)-3=-1 2(2)-3=1
domain (2,1) range (1,-1)
Please help.

Answer by stanbon(48558)

I need your help. I am not the greatest when it comes to solving functions.
x^2=1+y^2....is this a function or not
is this correct?
x^2-y^2=1
This is a hyperbola opening right and left.
It is not a function because an x value has more than one corresponding
y value.
y = [sqrt(x^2-1)] or y = -[sqrt(x^2-1)]
domain [1,infin) or (-inf,-1]
range: All Real Numbers
--------------
Cheers,
stan H.